OpenStudy (anonymous):

What volume of water will fit inside a cylindrical barrel that has a total surface area of 1885 in^2 and a radius of 10 inches? 6283 in^3 20 in^3 1885 in^3 18850 in^3 The perimeter of a rectangle is one more than the area of the rectangle. Find the area of the rectangle, if its width is 3 cm. 5 cm^2 16 cm^2 225 cm^2 15 cm^2

OpenStudy (anonymous):

Figure out the surface area of the cylinder, which will help you solve for h. After you get h, you need to find the volume which for a cylinder is just \[V = \pi r^2h\]

OpenStudy (anonymous):

the surface area is 2*pi*r^2

Directrix (directrix):

Total Surface Area of Cylinder = Lateral Area + Base Area -------------------------- Total Surface Area of a Cylinder = 2 pi r 2 + 2 pi r h In this problem, the total surface area is 1885 and r = 10 Total Area of a Cylinder = 2 pi r^ 2 + 2 pi r h

Directrix (directrix):

We will plug in for r = 10 and total surface area - 1885 and solve for h. Then, there will be enough info to get the volume. @kittymeow101

Directrix (directrix):

1885 = 2 pi *10^2 + 2 pi * 10* h

Directrix (directrix):

I take it we are to approximate for pi? @kittymeow101

OpenStudy (anonymous):

yes

Directrix (directrix):

Take this: 1885 = 2 pi *10^2 + 2 pi * 10* h What is 2 pi *10^2 =

Directrix (directrix):

2 times 3.14 times 100 =

OpenStudy (anonymous):

628

Directrix (directrix):

1885 = 2 pi *10^2 + 2 pi * 10* h 1885 = 628 + 2*pi*10*h What is this: 2*pi*10 = ? ---->

OpenStudy (anonymous):

=62.4

OpenStudy (anonymous):

62.8*

Directrix (directrix):

ok

Directrix (directrix):

1885 = 628 + 62.8*h 1885 - 628 = 62.8*h What is: 1885 - 628 = ?

OpenStudy (anonymous):

=1257

Directrix (directrix):

1885 - 628 = 62.8*h 1257 = 62.8*h What is 1257 divided by 62.8 = ?

Directrix (directrix):

You can put this in the Google search box and get the quotient: 1257/62.8=

OpenStudy (anonymous):

20.0

Directrix (directrix):

Yes.

OpenStudy (anonymous):

thats the final answer right

Directrix (directrix):

No.

Directrix (directrix):

We used the surface area formula and the given information to solve for h.

Directrix (directrix):

The question asks for volume and h is one of the variables in the formula so we had to solve for it before going on to volume. So, here we go.

Directrix (directrix):

The volume of a right circular cylinder is the area of its base times its height. V = π * r² * h

OpenStudy (anonymous):

v=pi*r^2*h

Directrix (directrix):

V = π * r² * h V = π * 10² * 20

Directrix (directrix):

V = π * 10² * 20 ------------------- What is 3.14 * 100 * 20 =

OpenStudy (anonymous):

=6280

Directrix (directrix):

I used an online calculator and did not approximate for pi myelf and got 6283 and some change. Let me run the 3.14 version through.

Directrix (directrix):

6280 --> what sort of units?

OpenStudy (anonymous):

i got 6280 bcs i got 3.14*100*20

Directrix (directrix):

Whoever wrote the questions and the ops did not use 3.14 for pi. That's okay.

Directrix (directrix):

I see the 6283 among the answer options.

OpenStudy (anonymous):

so the answer would be A?

Directrix (directrix):

That's what I think it has to be.

Directrix (directrix):

Next time, we will not approximate for pi unless told to do so.

Directrix (directrix):

Just let the online calculator use whatever number of digits of pi that it has been programmed to do.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

should we get started on the second question?

Directrix (directrix):

Let's look at the problem.

Directrix (directrix):